OpenStudy (mendicant_bias):

(Introductory Real Analysis) I'm trying to transform a given expression with absolute value signs into a version without it by applying the formal definition of the absolute value function (example below), but I'm running into trouble when I get more than one absolute value sign inside another.

OpenStudy (anonymous):

i could see how that would be annoying

OpenStudy (mendicant_bias):

(Sorry, just having trouble putting up the LaTeX, I'll try to put up a pic instead of the math maybe)

OpenStudy (mendicant_bias):

How do I do a piecewise function bracket again? I'm trying to do it, but am somewhere making a mistake.

OpenStudy (anonymous):

let me do one, you can copy it

OpenStudy (mendicant_bias):

I'm trying this: \begin{displaymath} f(x) = \left\{ \begin{array}{lr} 1 & : x \in \mathbb{Q}\\ 0 & : x \notin \mathbb{Q} \end{array} \right. \end{displaymath} to display the abs. value function, and it's not rendering

OpenStudy (mendicant_bias):

Sure, thanks.

OpenStudy (anonymous):

\[f(x) = |x+4| = \left\{\begin{array}{rcc} x + 4 & \text{if} & x \geq -4 \\ - x - 4& \text{if} & x < -4 \end{array} \right.\]

OpenStudy (mendicant_bias):

What the hell? I just typed it all up in the equation box, and when I entered, nothing was there. OS isn't playing nice with me atm.

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} 1 & \text{if} & x \in\mathbb{Q} \\ 0& \text{if} & x\notin\mathbb{Q} \end{array} \right.\]

OpenStudy (anonymous):

not really an absolute value equation though

OpenStudy (anonymous):

or function or whatever

OpenStudy (anonymous):

\[ f(x) = \left\{ \begin{array}{lr} 1 & : x \in \mathbb{Q}\\ 0 & : x \notin \mathbb{Q} \end{array} \right.\]

OpenStudy (anonymous):

that is what you wrote i just omitted the "displaymath" in favor of \[

ganeshie8 (ganeshie8):

i smell a continuity/discontinuity proof is on the way

OpenStudy (zzr0ck3r):

I think sense they are working with the definition of absolute, they will be doing stuff like triangle inequality, and results of it :)

OpenStudy (anonymous):

@Mendicant_Bias you know how to copy this right?

OpenStudy (anonymous):

oh i thought it would be something like \[f(x)=||x+3|-2|\]

OpenStudy (mendicant_bias):

...And now it is, nevermind, just tremendous lag/delay between it rendering.

OpenStudy (anonymous):

draw it!

OpenStudy (zzr0ck3r):

What are you trying to do with it, most of the time we don't use the definition.

OpenStudy (mendicant_bias):

I'm trying to change the expression to not have the absolute value signs, which I did by using (max) and (min). http://i.imgur.com/l7T5TkD.png

OpenStudy (anonymous):



OpenStudy (anonymous):

looks right to me

OpenStudy (mendicant_bias):

What I did was simplify the insides and then use max on this one, e.g. \[a+b+|a-b|=2 \ \text{max}(a,b)\]

OpenStudy (anonymous):

course the first one is \(2a\) and the second is \(2b\)

OpenStudy (mendicant_bias):

\[a+b+2x+|a-b|+|a+b-2c+|a-b||\] I'd in principle, like to do the same thing, but I don't know what will happen/how the absolute value inside another one will work, really.

OpenStudy (mendicant_bias):

Whoops, no x, that's c.

OpenStudy (anonymous):

yeah was afraid of those nested absolute values

ganeshie8 (ganeshie8):

\(a\gt b\) : \[a+b+2x+|a-b|+|a+b-2c+|a-b|| \\= a+b+2x+a-b + |a+b-2c+a-b|\\=2a+2x + |2a-2c|\]

ganeshie8 (ganeshie8):

you just follow the rules but what exactly is the actual problem

OpenStudy (mendicant_bias):

"Write the following expressions in equivalent forms not involving absolute values."

ganeshie8 (ganeshie8):

ahh that really going to require some clever thinking..

OpenStudy (mendicant_bias):

Do you have to do it recursively, then? e.g. are there four possible outcomes or something like that, or do you just take the same condition you first applied to the outside absolute value signs and apply it again? Yeah, and I've never actually seen or used the max/min(value, value, value...) notation, came up unexpectedly.

OpenStudy (mendicant_bias):

But yeah, x is actually c, typo

ganeshie8 (ganeshie8):

oh then its easy

ganeshie8 (ganeshie8):

\(a\gt b\) : \[a+b+2c+|a-b|+|a+b-2c+|a-b|| \\= a+b+2c+a-b + |a+b-2c+a-b|\\=2a+2c + |2a-2c|\\ = 2max(a,c)\]

OpenStudy (mendicant_bias):

Guessing you just apply the same logic of the last problem after the first abs. val sign is gone

OpenStudy (mendicant_bias):

Yeah

ganeshie8 (ganeshie8):

Yes looks its similar..

OpenStudy (mendicant_bias):

Alright, cool. Gonna play with this for a second, but I think I got it.

ganeshie8 (ganeshie8):

does this work ? \[a+b+2c+|a-b|+|a+b-2c+|a-b|| \\~\\ = 2max(a+b+|a-b|, ~2c)\\~\\ = 2max(2max(a,b), 2c) \]

OpenStudy (mendicant_bias):

Wow, this is kind of bizarre. I don't know if I just haven't noticed something like this in math yet, but I don't think I've ever noticed something that feels like an operation/changing something (e.g. a>b) and it influencing an expression in layers all at once without having to break them down to get to the "inside" layer. It's not really math, it's an assumption, but I'm so used to order of ops. problems that having something change like that seems strange. The answer is a little different, and a little more aesthetically pleasing, if that's a hint, heh

OpenStudy (mendicant_bias):

Think I got it, one sec. http://i.imgur.com/V1jKffw.png Now turning that into max/min things.

ganeshie8 (ganeshie8):

4max(a,b,c) ?

OpenStudy (mendicant_bias):

Yeah, it's just \[4 \ \text{max(a,b,c)}\] Yeah, hah

OpenStudy (mendicant_bias):

Has a ring to how it's written. Well, cool! Thanks so much for helping me figure this out! I definitely wouldn't have gotten it had I not asked...lol

OpenStudy (mendicant_bias):

I didn't know that max worked like that, awesome.

OpenStudy (mendicant_bias):

(I mean, I didn't know how max worked at all a couple of hours ago, heh)

ganeshie8 (ganeshie8):

they are usually messy the given expression was cooked up to work out smoothly haha

ganeshie8 (ganeshie8):

\[\color{blue}{a+b+|a-b|}+\color{red}{2c}+|\color{blue}{a+b+|a-b|}+\color{red}{2c}| \\~\\ = 2max(\color{blue}{a+b+|a-b|}, ~\color{red}{2c})\\~\\ = 2max(\color{blue}{2max(a,b)}, \color{red}{2c})\\~\\ =4max(\color{blue}{max(a,b)},c)\\~\\ =4max(a,b,c) \]